OpenStudy (hy123):

-5(sqrt)27+(sqrt)12 simplify

OpenStudy (hy123):

\[-5\sqrt{27}+\sqrt{12}\]

OpenStudy (anonymous):

27=3*.......

OpenStudy (hy123):

Is the answer \[-13\sqrt{3}\]?

OpenStudy (anonymous):

\[-15\sqrt 3 +2 \sqrt 3\] guess you're right!

OpenStudy (hy123):

thanks :D